OpenStudy (anonymous):

Help simplifying trigonometric expressions:

OpenStudy (anonymous):

\[(1-\sin ^2x)/(sinx-cscx)\]

OpenStudy (apoorvk):

write csc(x) as 1/sinx. then multiply like binomial by binomial (like we did in quadratics). post and verify!!!

OpenStudy (anonymous):

So like this? \[(1-\sin ^2x)/(\sin x-(1/\sin x))\] Should we convert the 1-sin^2x to cos^2x? Or just continue with \[(1-\sin^2x)((1/\sin x)-\sin x)\]

OpenStudy (phi):

out of curiosity, how did you get from (1−sin2x)/(sinx−(1/sinx)) to (1−sin2x)((1/sinx)−sinx) ?

OpenStudy (anonymous):

Ooh, there was a mistake in multiplying by the reciprocal of the denominator.

OpenStudy (phi):

How about re-writing 1/(sinx -1/sinx) as sin(x)/ ( sin^2 x - 1) (by multiplying by sinx/sinx)

OpenStudy (phi):

or even better -sin(x)/ ( 1-sin^2 x ) by multiplying through by -1/-1

OpenStudy (anonymous):

so the expression would simplify to \[-\sin x\]

OpenStudy (phi):

that is what I got.

OpenStudy (anonymous):

Thank you very much!